# Balancing Efficiency and Expressiveness: Subgraph GNNs with Walk-Based Centrality

### v1

@@ Abstract
We propose an expressive and efficient approach that combines the strengths of two prominent extensions of Graph Neural Networks (GNNs): Subgraph GNNs and Structural Encodings (SEs). Our approach leverages walk-based centrality measures, both as a powerful form of SE and also as a subgraph selection strategy for Subgraph GNNs. By drawing a connection to perturbation analysis, we highlight the effectiveness of centrality-based sampling, and show it significantly reduces the computational burden associated with Subgraph GNNs. Further, we combine our efficient Subgraph GNN with SEs derived from the calculated centrality and demonstrate this hybrid approach, dubbed HyMN, gains in discriminative power. HyMN effectively addresses the expressiveness limitations of Message Passing Neural Networks (MPNNs) while mitigating the computational costs of Subgraph GNNs. Through a series of experiments on synthetic and real-world tasks, we show it outperforms other subgraph sampling approaches while being competitive with full-bag Subgraph GNNs and other state-of-the-art approaches with a notably reduced runtime.
@@ 1 Introduction
Graph Neural Networks (GNNs) (Scarselli et al., 2009; Gori et al., 2005; Micheli, 2009) have achieved great success in learning tasks with graph-structured data. Typically, GNNs are based on the message passing paradigm (Gilmer et al., 2017), in which node features are aggregated over their local neighborhood recursively, resulting in architectures known as Message Passing Neural Networks (MPNNs). MPNNs have been shown to suffer from limited expressive power : they are bounded by $1$-WL (Xu et al., 2018; Geerts & Reutter, 2023; Weisfeiler & Leman, 1968; Morris et al., 2019; 2023) cannot count certain substructures (Chen et al., 2020; Bouritsas et al., 2022; Tahmasebi et al., 2020) or solve graph bi-connectivity tasks (Zhang et al., 2023b).
In order to improve expressive power, some works augment the node features with Structural Encodings (SE) or Positional Encodings (PE) (Bouritsas et al., 2022; Dwivedi et al., 2021; Fesser & Weber, 2024). These predefined features provide additional information on the graph topology, leading to an improvement on benchmarks compared to using the raw features (Rampášek et al., 2022). Recently, Subgraph GNNs (Bevilacqua et al., 2021; Frasca et al., 2022; Zhang & Li, 2021; Cotta et al., 2021) have been proposed to overcome some of the expressivity limitations of MPNNs. This approach preserves equivariance while relying less on feature engineering. First, a Subgraph GNN transforms a graph into a “bag of subgraphs” based on a specific selection policy. These subgraphs are then processed by an equivariant architecture and aggregated to make graph- or node-level predictions. One common approach to generate subgraphs, known as node-marking, is to mark a single node in the graph (Papp & Wattenhofer, 2022). In this case, each subgraph is then “tied” to a specific node in the original graph, and a shared MPNN generates a representation for each subgraph.
@@
In this work, we present an approach that drastically reduces the complexity of Subgraph GNNs while maintaining their competitive performance. We achieve this by identifying easy-to-compute structural features that unlock a parsimonious yet effective subgraph selection strategy, while providing complementary information enhancing the model’s expressiveness.
We first identify a family of node centrality measures (Estrada & Rodriguez-Velazquez, 2005; Benzi & Klymko, 2013; 2014) as easy-to-compute scores that: (i) recapitulate well the extent to which a subgraph alters the graph representation; (ii) correlate with relevant substructure counts. In light of this, we propose to leverage these measures to efficiently and effectively reduce the size of bags of subgraphs without requiring learning of additional components. Specifically, we prioritize selecting marked subgraphs associated with the top-ranking nodes according to these centrality measures and, in particular, to the Subgraph Centrality by Estrada & Rodriguez-Velazquez (2005). Then, we propose to additionally interpret Centrality scores as Structural Encodings (CSEs), and to utilize them to augment the node features of the selected subgraphs. This combination of approaches is justified from an expressiveness perspective, as we show neither of the two subsume the other. We demonstrate node marking of (selected) subgraphs separates graphs that are not separable by CSE-based feature augmentations and, vice-versa, that CSEs allow to separate pairs indistinguishable to our subsampled Subgraph GNNs.
The resulting method, dubbed HyMN (Hybrid Marking Network), is an approach whereby Subgraph GNNs and Structural Encodings work in tandem to effectively overcome the expressiveness limitations of MPNNs and unlock the application of Subgraph GNNs to larger graphs previously out of reach. Our approach is provably expressive, does not require feature engineering or learnable sampling components and, importantly, maintains a low computational cost. The practical value of our approach is confirmed by the positive results achieved over a series of experimental analyses conducted over synthetic and real-world benchmarks. We show HyMN outperforms other subgraph selection strategies, and performs on par or better than full-bag Subgraph GNNs by only sampling one or two subgraphs. Additionally, HyMN is competitive to (and sometimes better than) Graph Transformers and other state-of-the-art GNNs, while, as we show through extensive wall-clock timing analyses, featuring a more contained computational run-time.
@@
Overall, our results validate our method as a simple, expressive and efficient approach that competes with state-of-the-art GNNs with reduced empirirical run-times.
@@ 2 Related Work
Expressive power of MPNNs. The expressive power of MPNNs has become a central research topic since they were shown to be bounded by the 1-WL isomorphism test (Morris et al., 2019; Xu et al., 2018; Morris et al., 2023). This has led to approaches which aim to obtain different representations for non-isomorphic but 1-WL equivalent graphs. These include using random features (Sato et al., 2021; Abboud et al., 2020), higher-order message passing schemes (Bodnar et al., 2021b; a; Morris et al., 2019; 2020b) and equivariant models (Maron et al., 2018; Vignac et al., 2020). One of the most common approaches is to inject positional and structural encodings into the input layer (Bouritsas et al., 2022; Fesser & Weber, 2024; Dwivedi et al., 2021; Kreuzer et al., 2021) using Laplacian PEs (Dwivedi & Bresson, 2020; Kreuzer et al., 2021; Wang et al., 2022; Lim et al., 2023; 2024) or distance information (Ying et al., 2021; Li et al., 2020). A random-walk based encoding (RWSE) was proposed in (Dwivedi et al., 2021) and was shown to be able to distinguish regular graphs such as the 1-WL-equivalent CSL graphs (Murphy et al., 2019).
Subgraph GNNs. A recent line of work has proposed representing a graph as a collection of subgraphs obtained by a specific selection policy (Zhang & Li, 2021; Cotta et al., 2021; Papp & Wattenhofer, 2022; Bevilacqua et al., 2021; Zhao et al., 2022; Papp et al., 2021; Frasca et al., 2022; Qian et al., 2022; Huang et al., 2023; Zhang et al., 2023a). These approaches, jointly referred to as Subgraph GNNs, allow to overcome the expressivity limitations of MPNNs without introducing predefined encodings. A powerful, common selection policy termed node marking involves generating a subgraph per node by marking that node in the original graph, with no connectivity alterations. Although it can increase expressive power beyond 1-WL (Frasca et al., 2022; You et al., 2021), this approach has a high computational complexity as we need to consider and process $N$ different subgraphs, where $N$ is the number of nodes in the original graph.
@@ 3.2 Effective Node Marking
$h^{(l)}_{v}=\phi^{(l)}\big{(}h^{(l-1)}_{v}+{\epsilon}\sum_{u\in N(v)}h^{(l-1)}_{u}\big{)}\qquad y_{G}=\phi^{(L+1)}\big{(}\sum_{v\in G}h^{(L)}_{v}\big{)}$ (2)
where $\phi^{(l)}$’s are update functions and $\phi^{(L+1)}$ is a prediction layer (all are parameterized as MLPs). By applying results from GNN stability studies in (Chuang & Jegelka, 2022) we put forward the following observation (see details in Appendix D):
@@
where $K^{l}_{\phi}$ is the Lipschitz constant of MLP $\phi^{(l)},l=1\dots L+1$, $A$ is the adjacency matrix of graph $G$ and $\lambda_{l}\in\mathbb{R}^{+}$ is a layer-wise weighting scheme (dependent on ${\epsilon}$).
@@
We note that the above observation uncovers an intriguing alignment with walk-based centrality measures: the “most important” nodes associate with the largest (discounted) cumulative numbers of walks (compare Equations 1 with term $(\mathcal{A})$ in Equation 3). This leads us to direct our focus to walk-based centrality measures as promising candidates for selecting which node to mark. Considering we would like to sufficiently alter the graph representation beyond the one from a standard MPNN, Observation 1 indicates that, among all possible nodes, those associated with small cumulative numbers of walks will be poor marking candidates. We propose to summarise this form of information via walk-based centrality measures. Our strategy will be to rank nodes based on their centrality values, and mark the top-scoring ones. In the following, we empirically verify the validity of this approach, while deferring readers to Appendix D for extensions and additional considerations on the above analysis.
@@
High-centrality marking induces the largest perturbations. As a first experiment, we examine the extent to which node centralities recapitulate the amount of perturbation induced by marking their corresponding nodes. We consider the same setting discussed above: marking one node to transition from $G$ to $S_{v}$, where $S_{v}$ is obtained from $G$ by marking node $v$. For a centrality measure $c$, we consider three cases: $v$ attains the minimum of $c$ (i), the maximum of $c$ (ii), is randomly picked (iii). In each of these cases, we measured the distance $\|f(S_{v})-f(G)\|$ on $100$ graphs from two different real-world datasets from the popular TU suite: MUTAG and NCI1 (Morris et al., 2020a). Here, $f$ is an untrained $3$-layer GIN (Xu et al., 2018). Figure 1 shows results for the walk-based Subgraph Centrality (Estrada & Rodriguez-Velazquez, 2005), where horizontal lines indicate the average representation distance, and (i), (ii), (iii) are color-coded, resp., in green, blue, red.
@@
From the plots, it is clearly visible how marking nodes with the lowest centrality leads to the smallest change in graph representation. This result gives direct empirical validation to the upper-bound analysis and the consequential observation discussed above. In accordance with our claim, subgraphs associated with low centrality nodes can be interpreted as “redundant” w.r.t. the original input graph, or simply as poor marking candidates. Second, we note that marking nodes with the highest centrality induces the highest average perturbations, above random marking. This result is particularly relevant as it complements the above theoretical analysis: the walk-based upper-bound (Equation 3) only suggests, but does not necessarily entail, that high-centrality marking associates with the largest perturbations. Figure 1 shows that this occurs in practice on these datasets, further motivating our proposed sampling strategy. Results for other centrality measures are found in Appendix E; they indicate that high-centrality marking leads, in all cases, to larger perturbations than random marking, with the highest values attained by walk-based centrality measures (see Section 2).
@@
We randomly generate $100$ Erdös-Renyi (ER) graphs, each with $N=20$ nodes and wiring probability $p=0.3$. Similarly as above, we experiment with various centrality measures, by marking a single node $v$ which attains either the maximum or minimum centrality value, or is randomly picked. Again, we record the perturbation $\|f(S_{v})-f(G)\|$ given by the same architecture described above. On the same graphs, we count the number of various substructures, and evaluate the Pearson correlation between this value and the recorded perturbations. Results are reported in Table 1.
@@
Discussion. Overall our experiments indicate the following. First, high-centrality sampling appears to be a better approach than random sampling, especially when walk-based centrality measures are employed: on average, it selects marking candidates inducing the largest amount of perturbations over the original graph representations, and in a way that correlates with counts of relevant graph substructures. Second, the walk-based Subgraph Centrality stands out as a particularly promising candidate: it is efficient to precalculate this measure and sample subgraphs based on that, while, on average, it performed as the best one in the experiments discussed above.
In the following, we will focus on this measure in particular. Our strategy will consist of marking only the top-ranking $k$ nodes according to SC, for a small, fixed $k$. As we will show, this simple approach already delivers strong empirical performance. Note that more sophisticated sampling schemes could be designed based on extensions of the above perturbation analysis. These could consider more complex Subgraph GNN architectures (Frasca et al., 2022; Zhang et al., 2023a) or study the effect of multiple node markings, for which a deeper inquiry could take into account pair-wise scores beyond node-wise centrality measures [^2] For example, when selecting multiple nodes to mark, one could also account for the distance between them.. We defer these efforts to future work.
(a) MUTAG perturbation distances
(b) NCI1 perturbation distances
Figure 1: Plots showing the amount the graph representation using GIN is altered by adding an additional node-marked subgraph with (i) the highest centrality, (ii) the lowest centrality and (iii) a random marking. This is shown for both the MUTAG dataset 1(a) and for the NCI1 dataset 1(b).
Table 1: Pearson correlation between various substructure counts and amount of perturbation caused by different node-marking policies.
@@
Maximum Subgraph Centrality 0.947 0.956 0.958 0.972
Minimum Subgraph Centrality 0.644 0.643 0.634 0.698
@@
Maximum Degree Centrality 0.937 0.947 0.948 0.962
Maximum Closeness Centrality 0.935 0.937 0.946 0.957
Maximum Betweenness Centrality 0.803 0.816 0.821 0.845
4 Combining Subgraph GNNs with Structural Encodings

### v2

@@ Abstract
Subgraph GNNs have emerged as promising architectures that overcome the expressiveness limitations of Graph Neural Networks (GNNs) by processing bags of subgraphs. Despite their compelling empirical performance, these methods are afflicted by a high computational complexity: they process bags whose size grows linearly in the number of nodes, hindering their applicability to larger graphs. In this work, we propose an effective and easy-to-implement approach to dramatically alleviate the computational cost of Subgraph GNNs and unleash broader applications thereof. Our method, dubbed HyMN, leverages walk-based centrality measures to sample a small number of relevant subgraphs and drastically reduce the bag size. By drawing a connection to perturbation analysis, we highlight the strength of the proposed centrality-based subgraph sampling, and further prove that these walk-based centralities can be additionally used as Structural Encodings for improved discriminative power. A comprehensive set of experimental results demonstrates that HyMN provides an effective synthesis of expressiveness, efficiency, and downstream performance, unlocking the application of Subgraph GNNs to dramatically larger graphs. Not only does our method outperform more sophisticated subgraph sampling approaches, it is also competitive, and sometimes better, than other state-of-the-art approaches for a fraction of their runtime.
Machine Learning, ICML
@@ 1 Introduction
Graph Neural Networks (GNNs) (Scarselli et al., 2009; Gori et al., 2005; Micheli, 2009) have achieved great success in learning tasks with graph-structured data. Typically, GNNs are based on the message passing paradigm (Gilmer et al., 2017), in which node features are aggregated over their local neighborhood recursively, resulting in architectures known as Message Passing Neural Networks (MPNNs). MPNNs have been shown to suffer from limited expressive power: they are bounded by $1$-WL (Xu et al., 2018; Geerts & Reutter, 2023; Weisfeiler & Leman, 1968; Morris et al., 2019, 2023) cannot count certain substructures (Chen et al., 2020; Bouritsas et al., 2022; Tahmasebi et al., 2020) or solve graph bi-connectivity tasks (Zhang et al., 2023b).
Recently, Subgraph GNNs (Bevilacqua et al., 2021; Frasca et al., 2022; Zhang & Li, 2021; Cotta et al., 2021) have been proposed to overcome some of the expressivity limitations of MPNNs. This approach preserves equivariance while relying less on feature engineering. First, a Subgraph GNN transforms a graph into a “bag of subgraphs” based on a specific selection policy. These subgraphs are then processed by an equivariant architecture and aggregated to make graph- or node-level predictions. One common approach to generate subgraphs, known as node-marking, is to mark a single node in the graph (Papp & Wattenhofer, 2022). In this case, each subgraph is then “tied” to a specific node in the original graph, and a shared MPNN generates a representation for each subgraph.
@@
Summarizing, while Subgraph GNNs have demonstrated promising results on small molecular graphs, their computational requirements make them impractical for graphs larger than a few tens of nodes. This severe limitation needs to be addressed in an effective way to broaden the application of these models.
Our approach. In this work, we present an approach that drastically reduces the complexity of Subgraph GNNs while maintaining their competitive performance. We achieve this by identifying easy-to-compute structural features that unlock a simple yet effective subgraph selection strategy, while providing complementary information enhancing the model’s expressiveness.
We first identify a family of node centrality measures (Estrada & Rodriguez-Velazquez, 2005; Benzi & Klymko, 2013, 2014) as easy-to-compute scores that: (i) recapitulate well the extent to which a subgraph alters the graph representation; (ii) correlate with relevant substructure counts. In light of this, we propose to leverage these measures to efficiently and effectively reduce the size of bags of subgraphs without requiring additional learnable components. Specifically, we prioritize selecting marked subgraphs associated with the top-ranking nodes according to these centrality measures and, in particular, to the Subgraph Centrality by Estrada & Rodriguez-Velazquez (2005). Then, we propose to additionally interpret centrality scores as Structural Encodings (SEs) (Bouritsas et al., 2022; Dwivedi et al., 2021; Fesser & Weber, 2024), and to utilize them to augment the node features of the selected subgraphs. This combination of approaches is justified from an expressiveness perspective, as we show neither of the two subsume the other. We demonstrate node marking of (selected) subgraphs separates graphs that are not separable by Centrality-based SE (CSE) feature augmentations and, vice-versa, that CSEs allow to separate pairs indistinguishable to our subsampled Subgraph GNNs.
The resulting method, dubbed HyMN (Hybrid Marking Network), is an approach whereby Subgraph GNNs and Structural Encodings work in tandem to effectively overcome the expressiveness limitations of MPNNs and unlock the application of Subgraph GNNs to larger graphs previously out of reach. Our approach is provably expressive, does not require feature engineering or learnable sampling components, and, importantly, maintains a low computational cost. The practical value of our approach is confirmed by the strong results achieved over a series of experimental analyses conducted over synthetic and real-world benchmarks. We show HyMN outperforms other subgraph selection strategies, and performs on par or better than full-bag Subgraph GNNs by only sampling one or two subgraphs. Additionally, HyMN is competitive to (and sometimes better than) Graph Transformers and other state-of-the-art GNNs, while, as we show through extensive wall-clock timing analyses, features a dramatically reduced computational run-time.
@@
  4. 4.
    We show how our approach enables subgraph-methods to scale to substantially larger graphs than previously possible, even tens or hundreds of times larger.
Overall, our results validate our method as a simple, expressive and efficient approach that competes with state-of-the-art GNNs for a fraction of their empirirical run-times.
@@ 2 Related Work
Expressive power of MPNNs. The expressive power of MPNNs has become a central research topic since they were shown to be bounded by the 1-WL isomorphism test (Morris et al., 2019; Xu et al., 2018; Morris et al., 2023). This has led to approaches which aim to obtain different representations for non-isomorphic but 1-WL equivalent graphs. These include using random features (Sato et al., 2021; Abboud et al., 2021), higher-order message passing schemes (Bodnar et al., 2021b, a; Morris et al., 2019, 2020b) and equivariant models (Maron et al., 2018; Vignac et al., 2020). One of the most common approaches is to inject positional and structural encodings into the input layer (Bouritsas et al., 2022; Fesser & Weber, 2024; Dwivedi et al., 2021; Kreuzer et al., 2021) using Laplacian PEs (Dwivedi & Bresson, 2021; Kreuzer et al., 2021; Wang et al., 2022; Lim et al., 2023, 2024), distance information (Ying et al., 2021; Li et al., 2020), or random-walk encodings (RWSE) (Dwivedi et al., 2021).
Subgraph GNNs. A recent line of work has proposed representing a graph as a collection of subgraphs obtained by a specific selection policy (Zhang & Li, 2021; Cotta et al., 2021; Papp & Wattenhofer, 2022; Bevilacqua et al., 2021; Zhao et al., 2022; Papp et al., 2021; Frasca et al., 2022; Qian et al., 2022; Huang et al., 2023; Zhang et al., 2023a). These approaches, jointly referred to as Subgraph GNNs, allow to overcome the expressivity limitations of MPNNs without introducing predefined encodings. A powerful, common selection policy termed node marking involves generating a subgraph per node by marking that node in the original graph, with no connectivity alterations. Although it can increase expressive power beyond 1-WL (Frasca et al., 2022; You et al., 2021), this approach has a high complexity as it needs to consider and process $N$ different subgraphs, $N$ being the number of nodes in the original graph.
@@ 3.2 Effective Node Marking
$h^{(l)}_{v}\!=\!\phi^{(l)}\big{(}h^{(l-1)}_{v}+{\epsilon}\!\!\!\sum_{u\in N(v)}\!\!\!h^{(l-1)}_{u}\big{)},\ y_{G}\!=\!\phi^{(L+1)}\big{(}\sum_{v\in G}h^{(L)}_{v}\big{)}$ (2)
where $\phi^{(l)}$’s are update functions and $\phi^{(L+1)}$ is a prediction layer (all are parameterized as MLPs). By applying results from GNN stability studies in (Chuang & Jegelka, 2022) we put forward the following observation (see details in Appendix E):
@@
where $K^{l}_{\phi}$ is the Lipschitz constant of MLP $\phi^{(l)},l=1\dots L+1$, $A$ is the adjacency matrix of graph $G$ and $\lambda_{l}\in\mathbb{R}^{+}$ is an ${\epsilon}$-dependent layer-wise weighting scheme.
@@
We note that the above observation uncovers an intriguing alignment with walk-based centrality measures: the “most important” nodes associate with the largest (discounted) cumulative numbers of walks (compare Equations 1 with term $(\mathcal{A})$ in Equation 3). This leads us to direct our focus to walk-based centrality measures as promising candidates for selecting which node to mark. Considering we would like to sufficiently alter the graph representation beyond the one from a standard MPNN, Observation 1 indicates that, among all possible nodes, those associated with small cumulative numbers of walks will be poor marking candidates. We propose to summarise this form of information via walk-based centrality measures. Our strategy will be to rank nodes based on their centrality values, and mark the top-scoring ones. Next, we empirically verify the validity of this approach, and defer readers to Appendix E for extensions and more considerations on the above analysis.
@@
High-centrality marking induces the largest perturbations. As a first experiment, we examine the extent to which node centralities recapitulate the amount of perturbation induced by marking their corresponding nodes. We consider the same setting discussed above: marking one node to transition from $G$ to $S_{v}$, where $S_{v}$ is obtained from $G$ by marking node $v$. For a centrality measure $c$, we consider three cases: $v$ attains the minimum of $c$ (i), the maximum of $c$ (ii), is randomly picked (iii). In each of these cases, we measured the distance $\|f(S_{v})-f(G)\|$ on $100$ graphs from two different real-world datasets from the popular TU suite: MUTAG and NCI1 (Morris et al., 2020a). Here, $f$ is an untrained $3$-layer GIN (Xu et al., 2018). Figure 1 and Figure 7(a) show results for the walk-based Subgraph Centrality (Estrada & Rodriguez-Velazquez, 2005), where horizontal lines indicate the average representation distance, and (i), (ii), (iii) are color-coded, resp., in green, blue, red.
@@
From the plots, it is clearly visible how marking nodes with the lowest centrality leads to the smallest change in graph representation. This result gives direct empirical validation to the upper-bound analysis and the consequential observation discussed above. In accordance with our claim, subgraphs associated with low centrality nodes can be interpreted as poor marking candidates. Second, we note that marking nodes with the highest centrality induces the highest average perturbations, above random marking. This result is particularly relevant as it complements the above theoretical analysis: the walk-based upper-bound (Equation 3) only suggests, but does not necessarily entail, that high-centrality marking associates with the largest perturbations. Figure 1 and Figure 7(a) show that this occurs in practice on these datasets, further motivating our proposed sampling strategy. Results for other centrality measures are found in Appendix F; they indicate that high-centrality marking leads, in all cases, to larger perturbations than random marking, with the highest values attained by walk-based centrality measures (see Section 2).
@@
We randomly generate $100$ Erdös-Renyi (ER) graphs, each with $N=20$ nodes and wiring probability $p=0.3$. Similarly as above, we experiment with various centrality measures, by marking a single node $v$ which attains either the maximum or minimum centrality value, or is randomly picked. Again, we record the perturbation $\|f(S_{v})-f(G)\|$ given by the same architecture described above. On the same graphs, we count the number of various substructures, and evaluate the correlation between this value and the recorded perturbations. Results are reported in Table 1, expressed in terms of the Pearson correlation coefficient. Higher coefficients indicate substructure counts are better recapitulated by the perturbations induced by a certain marking strategy.
@@
Discussion. Overall our experiments indicate the following. First, high-centrality sampling shows to be a better approach than random sampling, especially when walk-based centrality measures are employed: on average, it selects marking candidates inducing the largest amount of perturbations over the original graph representations, and in a way that correlates with counts of relevant graph substructures. Second, the walk-based SC stands out as a particularly promising candidate: it is efficient to precalculate this measure and sample subgraphs based on that, while, on average, it performed as the best one in the experiments discussed above. In the following, we will focus on this measure in particular. Our strategy will consist of marking only the top-ranking $k$ nodes according to SC, for a small, fixed $k$. As we will show, this simple approach already delivers strong empirical performance. More sophisticated sampling schemes could be designed based on extensions of the above analyses. These could consider more complex Subgraph GNN architectures (Frasca et al., 2022; Zhang et al., 2023a) or study the effect of multiple node markings, for which a deeper inquiry could take into account pair-wise scores beyond node-wise centrality measures [^2] For example, when selecting multiple nodes to mark, one could also account for the distance between them.. We defer these efforts to future work.
Figure 1: Plot showing the amount the graph representation using GIN is altered in the MUTAG dataset by adding an additional node-marked subgraph with (i) the highest centrality, (ii) the lowest centrality and (iii) a random marking.
Table 1: Pearson correlation between substructure counts and perturbation caused by different node-marking policies.
@@
Max Subgraph Centrality 0.947 0.956 0.958 0.972
Min Subgraph Centrality 0.644 0.643 0.634 0.698
@@
Max Degree Centrality 0.937 0.947 0.948 0.962
Max Closeness Centrality 0.935 0.937 0.946 0.957
Max Betweenness Centrality 0.803 0.816 0.821 0.845
4 Combining Subgraph GNNs with SEs